# Electrons herald non-classical light

### v1

@@ Acknowledgements
Funding Information: All samples were fabricated in the Center of MicroNanoTechnology (CMi) at EPFL. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-15-1-0250 and was further supported by the Swiss National Science Foundation under grant agreements 185870 (Ambizione), 182103, and 176563 (BRIDGE). All experiments were carried out at the Göttingen UTEM Lab, which was funded by the Max Planck Society, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through 432680300/SFB 1456 (project C01) and the Gottfried Wilhelm Leibniz program, and the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101017720 (FET-Proactive EBEAM). Y.Y. acknowledges support from the EU H2020 research and innovation program under the Marie Sklodowska-Curie IF grant agreement No. 101033593 (SEPhIM).

### v2

@@ Acknowledgements
Funding Information: All samples were fabricated in the Center of MicroNanoTechnology (CMi) at EPFL. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-15-1-0250. All experiments were carried out at the Göttingen UTEM Lab, which was funded by the Max Planck Society, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through 432680300/SFB 1456 (project C01) and the Gottfried Wilhelm Leibniz program, and the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101017720 (FET-Proactive EBEAM). Y.Y. acknowledges support from the EU H2020 research and innovation program under the Marie Sklodowska-Curie IF grant agreement No. 101033593 (SEPhIM).